OpenStudy (anonymous):

Can someone please help me. I don't know what to do after: x+2/3=x+4/x->x^2+2x=3x+12. By the way I crossed multiplied to get x^2+2x=3x+12.

OpenStudy (anonymous):



OpenStudy (anonymous):

Ummm, you're assuming they're proportional, but I don't understand where you are getting that idea from.

OpenStudy (anonymous):

Do they form right triangles?

OpenStudy (anonymous):

If so, then we can use Pythagorean theorem to say \[ \sqrt{x^2-3^2} = \sqrt{(x+4)^2-(x+2)^2} \]

OpenStudy (anonymous):

Since x is the hypotenuse of the small triangle and x+4 is the hypotenuse of the big triangle.

OpenStudy (anonymous):

Squaring both sides gives us \[ x^2-3^2 = (x+4)^2-(x+2)^2 \]

OpenStudy (precal):